Blake and Matthew are partners who agree that Blake will receive 100,000 salary allowance and that any remaining income or loss will be shared equality. If Matthew’s capital account is credited for 2,000 as his share of the net income in a given period, how much net income did the partnership earn in that period
Answer #1
Share of net income credited to Mathew= $ 2000
Remainig income after salary allowance is shared equally
Hence,Remainig income after salary allowance=2000*2=$ 4000
Net income earned=Remaining income after salary allowance+ Salary allowance=4000+100000=$ 104000
I appreciate your ratings